an investigator invested a total of $1200 in two mutual funds. One fund earned a 6% profit while the other earned a 4% profit if the investors total profit was $54 how much was invested in each mutual fund.
Found 2 solutions by stanbon, amalm06:
Answer by stanbon(75887)

an investigator invested a total of $1200 in two mutual funds. One fund earned a 6% profit while the other earned a 4% profit if the investors total profit was $54 how much was invested in each mutual fund.
---------
Quantity Eq:: s + f = 1200
Interest Eq:: 6s+4f = 5400
-------------------------------------------
Modify for elimination::
4s + 4f = 4*1200
6s + 4f = 5400
----------------------
2s = 600
s = $300 (amt invested at 6%)
------
Solve for "f":
f = 1200-600 = $600 (amt invested at 4%)
--------------
Cheers,
Stan H.

Answer by amalm06(224)

Solve with one equation:
0.06x+(0.04)(1200-x)=54
0.06x+48-0.04x=54
0.02x=6
x=$300, at 6% (Answer)
$900-$300=$600, at 4% (Answer)